OpenStudy (anonymous):

how can 2.5% covert to 1.025?

myininaya (myininaya):

%=1/100. Replace the percent sign with 1/100 and multiply the 2.5

myininaya (myininaya):

Also, 2.5% doesn't convert to 1.025

myininaya (myininaya):

It does covert to 0.025 if that is what you meant

OpenStudy (anonymous):

2.5/100

OpenStudy (anonymous):

this is the deal

OpenStudy (anonymous):

I said that 2.5% in decimal is 0.025. My older family members have said 2.5% is 1.025.

OpenStudy (anonymous):

I have no idea how they got that number.??????????

myininaya (myininaya):

its impossible is why

myininaya (myininaya):

Don't believe everything you hear.

OpenStudy (anonymous):

pls see http://answers.yahoo.com/question/index?qid=20100721214016AAAmRXA LOOK4 An annual increase of 2.5% (or 0.025) means that each year, the salary is multiplied by 1.025. My supervisor agrees with the 1.025 too :(

OpenStudy (anonymous):

Finally I understood it. First, get the divisor : 2.5 / 100 = 0.025 0.025 + 1 = 1.025